OpenStudy (anonymous):

3. Segment AB is the diameter of circle M. The coordinates of A are (−4,3). The coordinates of M are (1,5). What are the coordinates of B? (1) (6,7) (3) (−3,8) (2) (5,8) (4) (−5,2) how do i show work???

OpenStudy (kainui):

First off, draw a picture and label the points and label B as (x,y) and solve with the pythagorean theorem since the distance between two points is the hypotenuse of a right triangle!

OpenStudy (anonymous):

\[(x-a)^2 + (y-b)^2 = r^2\]where M(a,b) is the center of the circle. We know that\[r = \sqrt{(x_A -x_M)^2 - (y_A - y_M)^2}\]Point B must lie on the circle and the line through Point A and Point M. The equation of the line through \(\bar{AB}\) is \[y - y_A = m(x-x_A)\]where \(m = {y_A - y_M \over x_A - x_M}\) We have two equations and two unknowns. We can solve one for an expression for x, plug that expression for x into the other, come up with a number for y, plug y back into the expression for x to get a value for x.